Check all of the expressions whose answers have more then 2 significant figures. A.7.893+10.1 B 144 times 4.1 C122.0 -47.4 D 200 times 3.5
1
Expert's answer
2020-12-17T19:08:57-0500

All expressions have more than 2 s.fs except D.


A. 7.893 + 10.1 = 17.993 (5s.f)

B. 144 * 4.1 = 590.4 (4s.f)

C. 122.0 - 47.4 = 74.6 (3s.f)

D. 200 * 3.5 = 700 (1s.f)


USEFUL TIPS:

  1. nonzero digits are significant. In 43,751, all the digits are significant.
  2. Zeros which are zeros between significant digits, are significant. In 704, all the digits are significant.
  3. Zeros at the beginning of a decimal number less than 1, are not significant. In 0.00037, the first four digits are not significant. The zeros serve only to put the digits 3 and 7 in the correct positions. This number has two significant figures.
  4. Zeros at the end of a number, are significant only if the number has a decimal point. Thus, in 2,600, the two trailing zeros are not significant because the number is written without a decimal point; the number has two significant figures. However, in 2,600.00, all six digits are significant because the number has a decimal point.

REFERENCE: Expressing Numbers: Significant Figures”, section 1.5 from the book Introduction to Chemistry: General, Organic, and Biological (v. 1.0).